OpenStudy (anonymous):

Help, plz Two large and 1 small pumps can fill a swimming pool in 4 hours. One large and 3 small pumps can also fill the same swimming pool in 4 hours. How many hours will it take 4 large and 2 small pumps to fill the swimming pool.(We assume that all large pumps are similar and all small pumps are also similar.)

OpenStudy (anonymous):

@whydoihavetosignup1

OpenStudy (anonymous):

@Phebe

OpenStudy (anonymous):

From the first information given you know that: 2L + 1S = 1 P => 4H The question now is: 4L + 2S = 1P in how many hours ?

OpenStudy (youngsuperman):

@TTP

OpenStudy (anonymous):

Evaluate it for a second and tell me if you notice anything about these two equations: 2L + 1S 4L + 2S

OpenStudy (phebe):

tbh i rly dont like Mathematics lolzz but ima see what i can do

OpenStudy (anonymous):

It's okay, I got this.

OpenStudy (phebe):

okay XD

OpenStudy (whydoihavetosignup1):

Yeah let him get it :)

OpenStudy (anonymous):

It depends on Savannah if she wants to learn how to solve it or if she just wants a clear answer.

OpenStudy (whydoihavetosignup1):

Can't solve this one bae any other subject i can do

OpenStudy (whydoihavetosignup1):

Human calculator gets the job done

OpenStudy (anonymous):

R u good at history?

OpenStudy (whydoihavetosignup1):

Get this question done first

OpenStudy (anonymous):

Ok

OpenStudy (phebe):

im very good at history

OpenStudy (anonymous):

So again, can you notice anything particular about the first information given, that: * 2 large pumps with 1 small pump fill 1 swimming pool (in 4 hours) And what is asked of you, that * 4 large pumps with 2 small pumps fill 1 swimming pool (in how many hours?)

OpenStudy (anonymous):

I got the answer! Thanks

OpenStudy (anonymous):

You don't even need the second information (that 1 large pump with 3 small pumps fill the swimming pool in 4 hours)

OpenStudy (anonymous):

Did you understand the principle though ?

OpenStudy (anonymous):

Yea, thanks for ur help

OpenStudy (whydoihavetosignup1):

Solve for R and r the system of first two equations then substitute in the third and solve for T to find the time. T = 5/3 hours = 1 hour 40 minutes. us what i found :/

OpenStudy (anonymous):

What the hell, no.

OpenStudy (anonymous):

There are twice as many pumps.

OpenStudy (anonymous):

That's what I wanted her to notice.

OpenStudy (whydoihavetosignup1):

ohh....

OpenStudy (anonymous):

So let's say 2 workers can complete something in 1 hour. Twice as many workers would complete the same thing in...30 minutes of course.

OpenStudy (phebe):

lolzzz!!! he said WTH no

OpenStudy (phebe):

at least dude tried on tha coool i didnt get the Q either

OpenStudy (phebe):

but now i do

OpenStudy (anonymous):

So here, 2 large pumps + 1 small pump complete something (filling a swimming pool) in 4 hours.

OpenStudy (anonymous):

Therefore, twice as many pumps (4 large and 2 small) complete the same thing (filling the same swimming pool) in half the time (2 hours).

OpenStudy (whydoihavetosignup1):

You da real MVP

OpenStudy (anonymous):

It's not a normal equation man, no worries. I thought maybe putting it in the form of an equation would help Savannah understand.

OpenStudy (anonymous):

It did! Thanks